The length of a rectangle is 5 meters longer than its width. If the perimeter of the garden is 50 meters, find the dimensions of the garden.
Answer by rfer(16322)

2w+2l=P
2(x+5)+2x=50
2x+10+2x=50
4x=40
x=10
x+5=15